OpenStudy (anonymous):

What is the solution to the equation? 16+x-11=18-5 x=?

OpenStudy (anonymous):

\[16+x-11=18-5\]??

OpenStudy (anonymous):

Did I type it in wrong?

OpenStudy (anonymous):

i am just asking if that is it first off, what is \(18-5\) ?

OpenStudy (lanhikari22):

uhh, no. 16 + x - 11 = 13 so 5 + x = 13 x = 8

OpenStudy (anonymous):



OpenStudy (anonymous):

is that what you would like answered?

OpenStudy (lanhikari22):

unless you meant 16 + (x-11) = 18-5, but i dunno

OpenStudy (anonymous):

18-5 = 13

OpenStudy (anonymous):

next?

OpenStudy (lanhikari22):

i meant 16 + x - 11 = (18-5)x actually, my bad